What is the mass of 8.30x1024 molecules of NH3?




1
Expert's answer
2021-05-03T07:40:18-0400

The molar mass of NH3 is 17.03 g/mol

1 mole of anything contains Avogadro's number of particles (6.022 x 1023). Therefore,


"m(NH_3)=8.30\times10^{24}molecules\times\frac{1mol(NH_3)}{6.022\times10^{23}molecules(NH_3)}\times\frac{17.03g(NH_3)}{1mol(NH_3)}=235g"


Answer: 235 g